the velocity v and displacement r of  body are related as v2=kr, whrer k is constant. what will be the velocity after 1 second? (give that the displacement is zero at t=0)
 

Grade:11

2 Answers

noogler
8 years ago
v=2r........................................v2=kr  (diff it w.r.t. t)
                                            2va=kv......k=2a              a is acceleratin
                                            v2=2ar............(v2-u2=2as)
                                           r=1/2at2.........(s=ut+1/2at2)
                                          a=2r/t2  substitute in v2=2ar.........v2=4r2/t2
kousheen
6 years ago
v2=kr
we can write it as v=\sqrt{}kr
diffentiating both sides
du/dt = \sqrt{}k  ½ r-1/2 dr/dt
a=\sqrt{}k ½ r-1/2 . v =\sqrt{}k ½ r-1/2 . \sqrt{}k r1/2
a=k/2 r0
velocity after 1 sec-- v=u+at =0+k/2 r0 X 1
so v=k/2 r0